Parth (parthkohli):

@Limitless @KingGeorge this is where the LaTeX will go as it doesn't work on chat.

Parth (parthkohli):

\(\Large {1 \over n} \in \mathbb{R} :n \ne 0\)

Parth (parthkohli):

Is this one correct?

Parth (parthkohli):

Oh \(\cdots :n\ne 0,\infty\)

OpenStudy (kinggeorge):

Your first was correct since \(\infty\notin\mathbb{R}\).

Parth (parthkohli):

I see.

OpenStudy (anonymous):

Two ways to define a function: \[\begin{align} f(x)&=\text{insert your expression in terms of } x\\ f:&X \to Y\\ &x \mapsto y \end{align}\] The second is under the condition that \(x\) is an element of \(X\) and \(y\) is an element of \(Y.\) You mean to write \[f(x)=\frac{1}{x} \,\,\, \{x: x \in \mathbb{R} \wedge x \ne 0\}.\]

OpenStudy (anonymous):

@KingGeorge, \(\frac{1}{n}\in \mathbb{R}: n\ne 0\) is acceptable?

Parth (parthkohli):

I meant to write that 1/n would be a real number if and only if n is not equal to infinity or zero/

OpenStudy (anonymous):

You mean, then, \[\left(\frac{1}{n} \in \mathbb{R}\right) \Leftrightarrow \left((n \ne 0) \vee (n \ne \infty)\right)\] You may exclude the parentheses.

OpenStudy (kinggeorge):

correct @Limitless \(\infty\) is not contained in the real numbers.

OpenStudy (kinggeorge):

Your first post was correct, not the most recent one.

OpenStudy (anonymous):

Interesting point, though. \(\mathbb{R}^{*}\) does consist of \(\infty\). It's called the extended real numbers. They are pretty cool.

OpenStudy (anonymous):

@KingGeorge, I'm thinking I flubbed up the last part there. I think it should be \(\wedge\) instead of \(\vee.\) What do you think?

Parth (parthkohli):

I heard that \(\mathbb{R}^*\) was supposed to be the non-zero real numbers?

OpenStudy (kinggeorge):

It should be \(\wedge\), but really, it shouldn't be there since the second statement is meaningless anyways.

Parth (parthkohli):

The lag has started. Let's insert the LaTeX only. REST TO CHAT.

OpenStudy (anonymous):

Sorry, it should be \(\bar{\mathbb{R}}\) for the extended real numbers.

OpenStudy (anonymous):

Also, \(^{*}\mathbb{R}\) is the hyperreal numbers.

Parth (parthkohli):

\(\Large {1 \over x} \in \mathbb{R} : x \in \mathbb{R}^{*}\)

OpenStudy (anonymous):

\[\{x: x\in \mathbb{R} \wedge x \ne 0\} \Rightarrow \frac{1}{x} \in \mathbb{R}\]

OpenStudy (anonymous):

seems legit

OpenStudy (kinggeorge):

\[S:=\left\{\frac{1}{x} :x\in \mathbb{R}^*\right\}\]

OpenStudy (anonymous):

\(\Leftrightarrow, \Rightarrow, \Leftarrow.\) See how they are all (almost) the same length?\(\iff\) clashes horribly.

Parth (parthkohli):

\(\Longleftarrow, \Longrightarrow, \Longleftrightarrow \). See how they are (almost) the same length? \(\Leftrightarrow \) clashes horribly.

OpenStudy (anonymous):

\Leftrightarrow \Leftarrow \Rightarrow \Longleftarrow \Longrightarrow \Longleftrightarrow Why use more letters when you don't need to?